OpenStudy (anonymous):

What value for a makes this equation true? 6 • a = (6 • 30) + (6 • 2) A. 36 B. 32 C. 28 D. 26

sammixboo (sammixboo):

Well let's simplify the right side of the equarion, (6*30) + (6*2)

sammixboo (sammixboo):

First multiply what is in the parenthesis

sammixboo (sammixboo):

Then add the two products together

sammixboo (sammixboo):

Tell me what you get

OpenStudy (igreen):

30 + 2 = ?

OpenStudy (anonymous):

32

OpenStudy (igreen):

6 * a = (6 * 30) + (6 * 2) ^ ^ 30 + 2

OpenStudy (anonymous):

is for 30 +2

OpenStudy (igreen):

Yes..that's our answer..

OpenStudy (anonymous):

HELP ME PLEASE

OpenStudy (igreen):

To check: 6 * 32 = 192 6 * 30 = 180 6 * 2 = 12 180 + 12 = 192

OpenStudy (anonymous):

now for multiplying i dont know wich i dio first

sammixboo (sammixboo):

I totally forgot about that little trick

sammixboo (sammixboo):

Hehe

OpenStudy (anonymous):

so oooooooooooooo q=what

OpenStudy (anonymous):

guys what do i do

sammixboo (sammixboo):

32